# Optimal market-neutral currency trading on the cryptocurrency platform

### v2

@@ 5.4.4 Risk and Sharpe Ratio
Profit and risk are the two sides of the same coin. We have shown records that the lower the Lambda parameter, the higher the profit. However, we also need to avoid the possibility of capital loss. The way we measure the risk of an investment is Sharpe Ratio Sharpe (1966). Sharpe $\text{Sharpe Ratio}=S_{p}=\frac{E[R_{p}]-R_{f}}{\sigma}$, where the $R_{p}$ is the portfolio’s annualized return, $R_{f}$ is the annual risk-free rate [^3] We use 4% based on U.S. 10-Year Treasury Note at the time of writing and $\sigma$ is the annualized standard deviation of the portfolio returns.
@@
Full-Cycle Market Bull Market Bear Market
Buy-Hold $S_{p}$: 0.12, $\sigma$: 0.961 $S_{p}$: 1.69, $\sigma$: 1.313 $S_{p}$: -0.61, $\sigma$: 0.870
DM $S_{p}$: -0.41, $\sigma$: 0.082 $S_{p}$: 0.60, $\sigma$: 0.273 $S_{p}$: -1.51, $\sigma$: 0.019
OTT($\lambda=2$) $S_{p}$: 0.44, $\sigma$: 0.104 $S_{p}$: 2.11, $\sigma$: 0.070 $S_{p}$: -0.32, $\sigma$: 0.042
OTT($\lambda=1$) $S_{p}$: 0.68, $\sigma$: 0.193 $S_{p}$: 2.66, $\sigma$: 0.126 $S_{p}$: 0.16, $\sigma$: 0.082
OTT($\lambda=0.5$) $S_{p}$: 1.11, $\sigma$: 0.343 $S_{p}$: 4.34, $\sigma$: 0.174 $S_{p}$: 0.41, $\sigma$: 0.161
  • 1
    $S_{p}$ is Sharpe Ratio of investment portfolio calculated based on a Risk-Free annual rate of 4%, $\sigma$ is the standard deviation of investment portfolio;
  • 2
    The full-cycle market is from 15-Jan-2021 to 01-Oct-2022. The bull market is from 01-Jan-2021 to 01-Jan-2022. And the bear market is from 01-Jan-2022 to 01-Jan-2023.
  • 3
    The participating techniques are ”BH” for simple buy-and-hold; ”DM” for Distance Method; ”OCT” for our Optimal Trading technique, where $\lambda$ represents are risk control factor;
  • 4
    Portfolio USD Equivalent is a conversion of measuring the trend of multiple assets back to a unified index. We extract the exchange rate from the last trading day, and use the exchange rate to convert all assets back to USD through the full trading period;
The results from Table 9 give comparisons of different techniques. From the annualized standard deviation $\sigma$, we can see that Distance Method (DM)and Optimal Trading Technique (OTT)under all conditions have significantly lower volatility than Buy-Hold. DM is less volatile than OTT because it trades less. For OTT, there is an expected negative relation between $\lambda$ and $\sigma$. A greater value of $\lambda$ effectively reduces portfolio volatility. When measuring the overall portfolio performance with the Sharpe Ratio, lower $\lambda$ brings higher profitability, therefore, a higher Sharpe Ratio.
@@ 5.5 Strategy Indicators
To have an overview of the trading strategy, we also look into the trading metrics, not just the profitability and risk. For the consistency of measurement, one action of arbitrage is the whole process of position opening and closing. One action of trade is a one-time position opening or closing, which consists of two buy or sell orders. Therefore, the relationship can be presented in an equation: 1 Arbitrage = 2 Trades [Position Open/Close] = 4 orders [Buy/Sell].
Table 10 includes some common trading strategy indicators. It consists of three sets of experiments: Full-Cycle, Bull and Bear markets based on Section 5.4. The Full-Cycle market contains the trades from 15-Jan-2021 to 01-Oct-2022 (625 days); the bull market contains 01-Jan-2021 to 01 Jan-2022 (365 days); and the bear market contains the trades from 01-Jan-2022 to 01-Jan-2023 (365 days).
  • number of trades The total number of trades happened during the whole arbitrage period
  • % of winning The percentage of winning trades among all trades
  • % of losing The percentage of losing trades among all trades
  • % of long win The percentage of winning trades among all long trades
  • % of short win The percentage of winning trades among all short trades
  • Win/Loss ratio The ratio of winning trades against losing trades
  • Average loss The average USD loss among losing trades
  • Average Win The average USD win among winning trades
  • Largest loss The largest USD loss among losing trades
  • Largest win The largest USD win among winning trades
  • Avg holding hrs on average how many hours is an arbitrage open
We can clearly see that our technique conducts trades way more frequently than the Distance Method. It is mainly because our spread focuses on the deviation of forex movement anchored to ETH while the Distance Method concentrates on the distance. Our six experiments setup has a Win/Loss ratio of greater than 1 (Table 10), which formed the foundation of pair trading profitability. It is reasonable that the longer the interval, the less trade and, therefore, longer average holding hours. We are also glad to see that there is no particularly large value in the largest win or largest loss that could impact the overall revenue. It is a symbol of the stability of our arbitrage.
@@
Indicators Fullcycle Optimal Trading Technique Distance Method
@@
Indicators Bull Optimal Trading Technique Distance Method
@@
Indicators Bear Optimal Trading Technique Distance Method
@@
33footnotetext: Optimal Trading technique adopts the most lucrative Open_threshold/Close_Threshold for 1min, 5min and 60min trading interval (Table 3)33footnotetext: All the data are collected under LAMBDA as 1 and transaction cost as 0.1%
As for the comparison of the bull market and bear market, there is a significant decrease in the number of trades and a slight drop in the average profit (or loss). That is to say, although our strategy remains effective in a bear market, the quality and quantity of the arbitrage opportunities are impacted.
@@ 6 Conclusion
In this paper, we examined the feasibility and profitability of multivariate pair trading. We take fiat currencies quoted for ETH as the bucket of our pair. We construct the spread focusing on the anomaly in forex, because of the movement of forex coming from independent trades against cryptocurrencies. The trading threshold is a tunable parameter based on the standard deviation of spreads. We demonstrated a grid search result of profitability under untuned conditions. In addition to the conventional pair trading strategy, we also introduced an optimization process for the resource allocation that emerged from multivariate pair trading, which focused on the restrictions of ensuring market neutrality and utilizing most of the holding assets (without leverage).
@@
The results are encouraging. The profitability of the proposed Optimal Trading Technique (OTT)is on average 2.72 times [^4] The average calculation in the conclusion section is calculated from 1min, 5min, 60min sampled dataset for bull, bear and full-cycle market under 0.1% transaction fee more profitable compared to the baseline approach – Distance Method. This implies the adoption of trading spread is more effective than the distance in the volatile cryptocurrency market. Under the assumption of a 4% risk-free interest rate, the Sharpe Ratio of OTT still achieved a strong 1.17. For all the tested cases, OTT achieved Win/Loss greater than 1, which is the driving force of steady capital growth.
@@
Our bi-criteria optimization objective can be tuned with risk tolerance parameter LAMBDA ($\lambda$), which impacts the optimization process for how much risk a trader tolerates. For the full-cycle market, 5min sampled dataset, the high ($\lambda=2$), intermediate ($\lambda=1$) and low ($\lambda=0.5$) risk aversion parameter value have annualized returns of 7.74%, 15.49% and 37.74%, respectively. The corresponding Sharpe Ratio 0.44, 0.68 and 1.11 have a negative correlation with $\lambda$, which means the strong profitability compensates for the volatility brought by riskier investment choices.
@@
Our Optimal Trading technique produces positive returns under both bull and bear market trends. In the bull market of 2021, OTT conducted 10456 trades, while in the bear market of 2022, OTT only conducted 3306 trades. Apparently the quantity and quality of arbitrage opportunities are better in a bull market.

### v3

@@ 5.4.4 Risk and Sharpe Ratio
Profit and risk are two sides of the same coin. We have shown that the lower the $\lambda$ parameter, the higher the profit. Risk is measured with the Sharpe Ratio Sharpe (1966). The Sharpe Ratio $S_{p}$ is calculated as follows: $S_{p}=\frac{E[R_{p}]-R_{f}}{\sigma}$ where $R_{p}$ is the portfolio’s annualized return, $R_{f}$ is the annual risk-free rate [^4] We use 4% based on the U.S. 10-Year Treasury Note at the time of writing, and $\sigma$ is the annualized standard deviation of the portfolio returns.
@@
Strategy Full-Cycle Market Bull Market Bear Market
$S_{p}$ $(\sigma)$ $S_{p}$ $(\sigma)$ $S_{p}$ $(\sigma)$
Buy-Hold 0.12 (0.961) 1.69 (1.313) -0.61 (0.870)
DM -0.41 (0.082) 0.60 (0.273) -1.51 (0.019)
OTTλ=2 0.44 (0.104) 2.11 (0.070) -0.32 (0.042)
OTTλ=1 0.68 (0.193) 2.66 (0.126) 0.16 (0.082)
OTTλ=0.5 1.11 (0.343) 4.34 (0.174) 0.41 (0.161)
  • * Portfolio value is calculated using the method described in Section 5.3.1.
  • * $S_{p}$ is the Sharpe Ratio of the investment portfolio calculated based on a risk-free annual rate of 4%.
  • * $\sigma$ is the standard deviation of the investment portfolio.
  • * $\lambda$ represents the risk control factor.
  • * The full-cycle market is from 15-Jan-2021 to 01-Oct-2022.
  • * The bull market is from 01-Jan-2021 to 01-Jan-2022.
  • * The bear market is from 01-Jan-2022 to 01-Jan-2023.
  • * BH (buy-and-hold), DM (Distance Method), OTT (Optimal Trading Technique).
The results in Table 9 provide comparisons of different techniques. From the annualized standard deviation $\sigma$, we can see that Distance Method (DM)and Optimal Trading Technique (OTT)under all conditions have significantly lower volatility than Buy-Hold. DM is less volatile than OTT because it trades less. For OTT, there is an expected negative relation between $\lambda$ and $\sigma$. A greater value of $\lambda$ effectively reduces portfolio volatility. When measuring the overall portfolio performance with the Sharpe Ratio, lower $\lambda$ brings higher profitability, and therefore, a higher Sharpe Ratio.
@@ 5.5 Strategy Indicators
To have an overview of the trading strategy, we also look into the trading metrics, not just the profitability and risk. One action of trade is a one-time position opening or closing, which consists of two buy or sell orders. Therefore, the relationship can be presented in an equation: 1 Arbitrage = 2 Trades [position open/close] = 4 Orders [buy/sell].
Table 10 includes some common trading strategy indicators. It consists of three sets of experiments: Full-Cycle, Bull, and Bear markets based on Section 5.4.
  • number of trades: The total number of trades happened during the whole arbitrage period
  • % winning: The percentage of winning trades among all trades
  • % losing: The percentage of losing trades among all trades
  • % long wins: The percentage of winning trades among all long trades
  • % short wins: The percentage of winning trades among all short trades
  • Win/Loss ratio: The ratio of winning trades against losing trades
  • Average loss: The average USD loss among losing trades
  • Average Win: The average USD win among winning trades
  • Largest loss: The largest USD loss among losing trades
  • Largest win: The largest USD win among winning trades
  • Avg holding hrs: on average how many hours is an arbitrage open
We can clearly see that our technique conducts trades much more frequently than the Distance Method. This is mainly because our spread focuses on the deviation of forex movements anchored to ETH, while the Distance Method concentrates on the distance. Our six experimental setups have a Win/Loss ratio greater than 1 (Table 10), forming the foundation of pair trading profitability. It is reasonable that the longer the interval, the fewer trades occur, resulting in longer average holding hours. We are also pleased to see that there are no particularly large values in the largest win or largest loss that could impact the overall revenue. This indicates the stability of our arbitrage strategy.
@@
Indicators (Fullcycle) Optimal Trading Technique Distance Method
@@
Indicators (Bull) Optimal Trading Technique Distance Method
@@
Indicators (Bear) Optimal Trading Technique Distance Method
@@
44footnotetext: Optimal Trading technique adopts the most lucrative Open_threshold/Close_Threshold for 1min, 5min and 60min trading interval (Table 4)44footnotetext: All the data are collected under $\lambda$ as 1 and transaction cost as 0.1%
As for the comparison between the bull market and bear market, there is a significant decrease in the number of trades and a slight drop in the average profit (or loss) in the bear market. This indicates that although our strategy remains effective in a bear market, the quality and quantity of arbitrage opportunities are impacted.
@@ 6 Conclusion
In this paper, we examined the feasibility and profitability of multivariate pair trading. We used fiat currencies quoted for ETH as the bucket of our pairs. We constructed the spread focusing on anomalies in forex movements due to independent trades against cryptocurrencies. The trading threshold is a tunable parameter based on the standard deviation of spreads. We demonstrated a grid search result of profitability under untuned conditions. In addition to the conventional pair trading strategy, we also introduced an optimization process for resource allocation in multivariate pair trading, ensuring market neutrality and utilizing most of the holding assets without leverage.
@@
The results are encouraging. The profitability of the proposed Optimal Trading Technique (OTT)is on average 2.72 times [^5] The average calculation in the conclusion section is based on 1min, 5min, and 60min sampled datasets for bull, bear, and full-cycle markets under a 0.1% transaction fee more profitable compared to the baseline approach—Distance Method. This implies that trading spread is more effective than distance in the volatile cryptocurrency market. Under the assumption of a 4% risk-free interest rate, the Sharpe Ratio of OTT achieved is a strong 1.17. For all the tested cases, OTT achieved a Win/Loss ratio greater than 1, which drives steady capital growth.
@@
Our bi-criteria optimization objective can be tuned with the risk tolerance parameter $\lambda$, which impacts the optimization process for how much risk a trader tolerates. For the full-cycle market, 5min sampled dataset, the high ($\lambda=2$), intermediate ($\lambda=1$), and low ($\lambda=0.5$) risk aversion parameter values have annualized returns of 7.74%, 15.49%, and 37.74%, respectively. The corresponding Sharpe Ratios of 0.44, 0.68, and 1.11 show a negative correlation with $\lambda$, indicating that strong profitability compensates for the volatility brought by riskier investment choices.
@@
Our Optimal Trading Technique produces positive returns under both bull and bear market trends. In the bull market of 2021, OTT conducted 10,456 trades, while in the bear market of 2022, OTT conducted only 3,306 trades. Clearly, the quantity and quality of arbitrage opportunities are better in a bull market.
Despite the promising results, several limitations should be noted. Cryptocurrency carries significant risks for investors, as highlighted by several sources. For instance, Swaminathan Anklesaria Aiyar (2022) likens the cryptocurrency market to the Dutch Tulip Bulb Bubble of the 17th century, where tulip bulb prices soared to extraordinary levels before collapsing dramatically. Additionally, many have documented the risks associated with cryptocurrency investments (Kerr et al. 2023; Angerer et al. 2021; Pakhnenko et al. 2023; LEE et al. 2018; Jing Pan 2023). We use cryptocurrency merely as an intermediate for pair trading fiat currencies due to its liquidity, 24/7 trading time, easy access to multiple fiat currencies in one market, and low barrier for data access. We do not promote investment in cryptocurrencies. Additionally, we selected two years of high-frequency data as our test data source, covering a full-cycle market, including a bull-run and a bear-run. However, this timeframe coincides with the COVID-19 period, during which fiat currencies were also volatile, contributing to spread movement. Profitability might differ in less volatile market conditions. Future work could extend the technique to model the spread from the perspective of the Ornstein–Uhlenbeck process to deepen the understanding of the market stochastic process.